# Optimal Compressed Sensing for Image Reconstruction with Diffusion Probabilistic Models

### v2

@@ E.2 CNN denoiser
We performed empirical experiments using UNet architecture. Our UNet networks contain 3 encoder blocks, one mid-level block, and 3 decoder blocks (Ronneberger et al., 2015). (For MNIST images we reduced the encoder and decoder blocks to 2, since they are smaller images.) Each block consists of 2 convolutional layers followed by a ReLU non-linearity and bias-free batch-normalization. Each encoder block is followed by a $2\times 2$ spacial down-sampling and a 2 fold increase in the number of channels. Each decoder block is followed by a $2\times 2$ spacial upsampling and a 2 fold reduction of channels. The total number of parameters is $7.6m$. All the denoisers are “bias-free”: we remove all additive constants from convolution and batch-normalization operations (i.e., the batch normalization does not subtract the mean). This facilitates universality (denoisers can operate at all noise levels) see Mohan et al. (2019).
We follow the training procedure described in Mohan et al. (2019), minimizing the mean squared error in denoising images corrupted by i.i.d. Gaussian noise with standard deviations drawn from the range $[0,1]$ (relative to image intensity range $[0,1]$). Training is carried out on batches of size $512$, for $1000$ epochs. Note that all denoisers are universal and blind: they are trained to handle a range of noise, and the noise level is not provided as input to the denoiser. These properties are exploited by the sampling algorithm, which can operate without manual specification of the step size schedule (Kadkhodaie & Simoncelli, 2020). This method produces high-quality results in generative sampling, as well as sampling conditioned on linear measurements (Kadkhodaie & Simoncelli, 2021). To train each denoiser, 4 NVIDIA A100 GPU were used. The total training time for each denoiser was around 10 hours.
@@ E.3 MMSE estimate
We construct the linear inverse estimate by averaging multiple samples obtained using Algorithm 2. The individual samples are shown in Figure 2 and Supplementary Figure 2 and Figure 3. Averaging multiple samples achieves a lower MSE by approximating the posterior mean, which we use when reporting performance (PSNR values). The number of samples $n$ is set to $n=2$ when optimizing the linear measurement to limit memory footprint, and $n=16$ when reporting performance on test data.
@@ E.4 Optimizing linear measurement
We search for the optimal linear measurement by performing stochastic gradient descent in the space of orthogonal matrices as described in the main text. The linear inverse procedure Algorithm 2 is end-to-end differentiable, and thus optimization of the measurement matrix can be done directly in PyTorch by taking derivatives of reconstruction loss with respect to the matrix parameterization. In all cases, we used the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$ with exponential decay of $0.90$. We used a batch size of $64$, and training was run for $16$ epochs. The optimization was done on a single node in a GPU cluster, with 4 NVIDIA A100 GPUs. The optimization for a single OLM requires a few hours for the MNIST dataset, and about 24 hours for the CelebA dataset.
@@ Appendix F Supplementary Figures and Tables
Supplementary Figure 1: Denoiser implicit prior, $\hat{x}(y)-y=\sigma^{2}\nabla\log p_{\sigma}(y)$, depicted as a vector field, for the three different data distributionsf Figure 1. Each vector indicates the direction in which the probability of the noisy density increases most rapidly (see Eq. 1).
@@
METHODS Bora et al. (2017) Wu et al. (2019) Our Method
@@
Table 1: Compare reconstruction performance to different compressed sensing methods. We compare the per-image MSE (lower is better) of our method with previous methods for compressed sensing using deep neural networks (Bora et al., 2017; Wu et al., 2019). Methods with symbol “*” use measurement function that are explicitly optimized; The symbol “/ ” indicates the results are not available, as Wu et al. (2019) reported performance only for small $k$; The symbol $\sim$” indicates the MSE are estimated from images of different size (64 x 64) assuming a constant per-pixel error. In addition, $\pm$ indicates standard error of the mean. Bold numbers highlight the best result for linear measurement.
We compare the per-image MSE of our method with previously proposed methods for compressed sensing using deep neural networks (Bora et al., 2017; Wu et al., 2019). Bora et al. (2017) used both VAE and GAN as image priors to obtain the linear inverse solution. Their methods outperform standard method for compressed sensing using LASSO (sparse prior) in the wavelet domain. They did not attempt to optimize the measurement matrix, and as a result were vastly outperformed by our method.
Wu et al. (2019) optimized both linear and nonlinear measurements using the restricted isometry property as an objective, but only for very small $k$. We found that on MNIST, our OLM is able to match the performance of even the nonlinear measurement function. On CelebA dataset, for $k=50$, we obtained slightly worse performance for the OLMs compared to the learned linear measurements in Wu et al. (2019). However, we may have underestimated the error in Wu et al. (2019), as we estimated MSE based on images of a smaller size (64 x 64) from their results but assumed a constant per-pixel error. Furthermore, our method is most effective for $k$ at around $10\%$ of the total number of pixels. Thus, we expect our advantage to improve further for larger $k$. Lastly, our method has the distinct feature that we can produce individual high-probability samples, in addition to the MMSE solution.

### v3

@@ E.2 CNN denoiser
We performed empirical experiments using UNet architecture. Our UNet networks contain 3 encoder blocks, one mid-level block, and 3 decoder blocks (Ronneberger et al., 2015). (For MNIST images we reduced the encoder and decoder blocks to 2, since they are smaller images.) Each block consists of 2 convolutional layers followed by a ReLU non-linearity and bias-free batch-normalization. Each encoder block is followed by a $2\times 2$ spacial down-sampling and a 2 fold increase in the number of channels. Each decoder block is followed by a $2\times 2$ spacial upsampling and a 2 fold reduction of channels. The total number of parameters is $7.6m$. All the denoisers are “bias-free”: we remove all additive constants from convolution and batch-normalization operations (i.e., the batch normalization does not subtract the mean). This facilitates universality (denoisers can operate at all noise levels) – see Mohan et al. (2019).
We follow the training procedure described in Mohan et al. (2019), minimizing the mean squared error in denoising images corrupted by i.i.d. Gaussian noise with standard deviations drawn from the range $[0,1]$ (relative to image intensity range $[0,1]$). Training is carried out on batches of size $512$, for $1000$ epochs. Note that all denoisers are universal and blind: they are trained to handle a range of noise, and the noise level is not provided as input to the denoiser. These properties are exploited by the sampling algorithm, which can operate without manual specification of the step size schedule (Kadkhodaie & Simoncelli, 2020). This method produces high-quality results in generative sampling, as well as sampling conditioned on linear measurements (Kadkhodaie & Simoncelli, 2021). To train each denoiser, 4 NVIDIA A100 GPU were used. The total training time for each denoiser was approximately 10 hours.
@@ E.3 MMSE estimate
We construct the linear inverse estimate by averaging multiple samples obtained using Algorithm 2. The individual samples are shown in Figure 2 and Supplementary Figure 3 and Figure 4. Averaging multiple samples achieves a lower MSE by approximating the posterior mean, which we use when reporting performance (PSNR values). The number of samples $n$ is set to $n=2$ when optimizing the linear measurement to optimize performance, as we found empirically that using more samples for the MMSE estimates during optimization does not improve performance during testing. For reporting performance on test data, we used $n=16$.
@@ E.4 Optimizing linear measurement
We search for the optimal linear measurement by performing stochastic gradient descent in the space of orthogonal matrices as described in the main text. The linear inverse procedure Algorithm 2 is end-to-end differentiable, and thus optimization of the measurement matrix can be done directly in PyTorch by taking derivatives of reconstruction loss with respect to the matrix parameterization. In all cases, we used the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-4}$ with exponential decay of $0.90$. We used a batch size of $64$, and training was run for $16$ epochs. The optimization was done on a single node in a GPU cluster, with 4 NVIDIA A100 GPUs. The optimization for a single OLM requires a few hours for the MNIST dataset, and about 24 hours for the CelebA dataset. The computational cost is primarily due to backpropagation through the diffusion model sampling procedure, which involves iterative applications of the U-Net denoiser.
@@ E.5 Subspace distance
E.6 Sequential OLM
The algorithm presented in Section 2.4 learns an OLM independently for each value of $k$. Empirically, we observe that the linear measurement subspace of smaller $k$ is typically not fully contained in that of larger $k$. In this section, we modify the algorithm into a greedy version, where measurement vectors are added one at a time, each conditioned on the previous measurement matrix. This results in a sequentially ordered set of OLMs, analogous to the structure of PCA.
The algorithm works as follows: Given a measurement matrix $M_{k}\in R^{d\times k}$, we find a set of vectors $U=\{u_{1},u_{2},...,u_{d-k}\}$ as a basis of the null space of $M_{k}$. The new measurement vector $m^{*}$ is parameterized as $m^{*}=U\lambda$, where $\lambda$ is a unit vector. The next measurement matrix $M^{*}\in R^{d\times(k+1)}$ is obtained by concatenating $M_{k}$ with $m^{*}$ as $M^{*}=[M_{k}\ m^{*}]$. The objective function in Equation 8 can be written with respect to $\lambda$ as:
$\mathcal{L}(\lambda)=\frac{1}{N}\sum_{i=1}^{N}||h(M^{*}(\lambda)^{T}x_{i};\ M^{*}(\lambda))-x_{i}||_{2}^{2}.$ (17)
To obtain a complete set of sequential OLM (denoted as $\text{OLM}_{\text{seq}}$), we start with $k_{0}=1$, and run the greedy algorithm iteratively, increasing the number of measurements $k$ by one at each step.
@@ Appendix F Supplementary Figures and Tables
Supplementary Figure 2: Denoiser implicit prior, $\hat{x}(y)-y=\sigma^{2}\nabla\log p_{\sigma}(y)$, depicted as a vector field, for the three different data distributions of Figure 1. Each vector indicates the direction in which the probability of the noisy density increases most rapidly (see Eq. 1).
@@
METHODS Bora et al. (2017) Wu et al. (2019b) Our Method
@@
Table 1: Compare reconstruction performance to different compressed sensing methods. We compare the per-image MSE (lower is better) of our method with previous methods for compressed sensing using deep neural networks (Bora et al., 2017; Wu et al., 2019b). Methods with symbol “*” use measurement function that are explicitly optimized; The symbol “/ ” indicates the results are not available, as Wu et al. (2019b) reported performance only for small $k$; The symbol $\sim$” indicates the MSE are estimated from images of different size (64 x 64) assuming a constant per-pixel error. In addition, $\pm$ indicates standard error of the mean. Bold numbers highlight the best result for linear measurement.
We compare the per-image MSE of our method with previously proposed methods for compressed sensing using deep neural networks (Bora et al., 2017; Wu et al., 2019b). Bora et al. (2017) used both VAE and GAN as image priors to obtain the linear inverse solution. Their methods outperform standard method for compressed sensing using LASSO (sparse prior) in the wavelet domain. They did not attempt to optimize the measurement matrix, and as a result were vastly outperformed by our method.
Wu et al. (2019b) optimized both linear and nonlinear measurements using the restricted isometry property as an objective, but only for very small $k$. We found that on MNIST, our OLM is able to match the performance of even the nonlinear measurement function. On CelebA dataset, for $k=50$, we obtained slightly worse performance for the OLMs compared to the learned linear measurements in Wu et al. (2019b). However, we may have underestimated the error in Wu et al. (2019b), as we estimated MSE based on images of a smaller size (64 x 64) from their results but assumed a constant per-pixel error. Furthermore, our method is most effective for $k$ at around $10\%$ of the total number of pixels. Thus, we expect our advantage to improve further for larger $k$. Lastly, our method has the distinct feature that we can produce individual high-probability samples, in addition to the MMSE solution.